# Coprime Bivariate Bicycle Codes and Their Layouts on Cold Atoms

### v5

@@ 1 Introduction
In contrast, research has demonstrated the existence of “good” qLDPC codes [13] indicating that qLDPC codes can have $k$ and $d$ scaling linearly as $n$ grows, i.e. codes with parameters $\llbracket n,k=\Theta(n),d=\Theta(n)\rrbracket$. However, having asymptotically good LDPC codes does not necessarily mean having better parameters than codes designed for short to medium lengths. Moreover, the structure of codes can greatly influence the design of hardware and vice versa. Therefore, for practical purposes, people started to seek qLDPC codes that are easy to implement on hardware and have good finite-length performance. For example, Panteleev and Kalachev [12] proposed generalized bicycle (GB) codes along with the BP-OSD decoder that focus on medium-length performance. It is noteworthy that they also discovered the $\llbracket 126,12,10\rrbracket$ GB code, which has the same parameters as the code we will analyze later. Another similar qLDPC code, the BB code [3], has received much attention because they have high thresholds, toric layout, and can be embedded on two planes. Benefiting from the quasi-cyclic and two-thickness properties of BB codes, recent papers have established the feasibility of BB codes on different architectures, including cold atom arrays [19, 14], trapped ion[9], and superconducting [1]. Due to the high connectivity of qLDPC codes, they are considered hard to implement on a superconducting-based platform unless a multi-layer architecture is used [1]. Benefiting from their all-to-all connectivity, cold atom array and trapped ion systems have certain advantages in implementing qLDPC codes. In particular, Viszlai et al. [19] showed that the quasi-cyclic structure of BB codes can be easily implemented using 2D atom array acousto-optic deflectors (AOD) for atom movement.
Constructing BB codes is, nonetheless, a time-consuming process as one has to search for combinations of polynomials to construct a code. Moreover, the parameters of codes can not be guaranteed. This challenge motivated us to develop an algorithm that accelerates the search for good BB codes and even constructs codes with pre-determined parameters. Our work differentiates itself from these approaches by introducing a general and efficient algorithm for searching BB codes in the form described in Eq. (5), which is also the form of codes used by related works. Furthermore, we proposed a novel algorithm that allows us to search for a subclass of BB codes with the desired dimensions, which we call coprime-BB codes. Coprime-BB codes generalize the form in Eq. (5) by allowing mixed terms but also have a restriction of using coprime circulant matrix sizes, making them unattainable through searches limited to the polynomial form of this equation. Thus, coprime BBs provide more flexibility for different scenarios on top of the originally proposed BB codes.
@@ 3.1 Code equivalence
share the same parameters, allowing us to search within only one class of these codes.
@@ 6 Related Work
Recently, various constructions based on qLDPC codes have been proposed to achieve different objectives. Koukoulekidis et al. [10] proposed algebraic extensions to expand a small GB code into a family of larger GB codes by selecting a sequence of expansion factors. The authors also introduced scalable codes that embed the original short codes into extended codes, enhancing scalability in superconducting architectures. Voss et al. [20] expanded the concept of BB codes by introducing an additional type of indeterminate variable, leading to the creation of trivariate bicycle (TB) codes. These new codes reduce the weight of stabilizers from 6 to 4-5, making them more practical for hardware implementation. However, it is important to note that some of these codes exhibit a lower rate or distance compared to BB codes of similar length. Eberhardt et al. [6] investigated the algebraic structure of BB codes and uncovered certain symmetry properties. These properties allow for the explicit construction of logical operators and certain fault-tolerant gates for BB codes. Shaw et al. [16] proposed a “morphing” circuit design for syndrome extraction on BB codes. The proposed circuit design has only six rounds of CNOT gates instead of seven [3]. By applying the proposed circuit, the authors discovered a new family of BB codes, including codes that have the same $\llbracket n,k,d\rrbracket$ parameters as [3]. The authors also provide a sufficient condition for the circuit to be applied to the other two-block group algebra (2BGA) codes.

### v6

@@ 1 Introduction
In contrast, research has demonstrated the existence of “good” qLDPC codes [3] indicating that qLDPC codes can have $k$ and $d$ scaling linearly as $n$ grows, i.e. codes with parameters $\llbracket n,k=\Theta(n),d=\Theta(n)\rrbracket$. However, having asymptotically good LDPC codes does not necessarily mean having better parameters than codes designed for short to medium lengths. Moreover, the structure of codes can greatly influence the design of hardware and vice versa. Therefore, for practical purposes, people started to seek qLDPC codes that are easy to implement on hardware and have good finite-length performance. For example, Panteleev and Kalachev [7] proposed generalized bicycle (GB) codes along with the BP-OSD decoder that focus on medium-length performance. It is noteworthy that they also discovered the $\llbracket 126,12,10\rrbracket$ GB code, which has the same parameters as the code we will analyze later. Another similar qLDPC code, the BB code [5], has received much attention because they have high thresholds, toric layout, and can be embedded on two planes. Benefiting from the quasi-cyclic and two-thickness properties of BB codes, recent papers have established the feasibility of BB codes on different architectures, including cold atom arrays [8, 9], trapped ion[10], and superconducting [11, 5]. Due to the high connectivity of qLDPC codes, they are considered hard to implement on a superconducting-based platform unless a multi-layer architecture is used [11]. Benefiting from their all-to-all connectivity, cold atom array and trapped ion systems have certain advantages in implementing qLDPC codes. In particular, Viszlai et al. [8] showed that the quasi-cyclic structure of BB codes can be easily implemented using 2D atom array acousto-optic deflectors (AOD) for atom movement.
Constructing BB codes is, nonetheless, a time-consuming process as one has to search for combinations of polynomials to construct a code. Moreover, the parameters of codes can not be guaranteed. This challenge motivated us to develop an algorithm that accelerates the search for good BB codes and even constructs codes with pre-determined parameters. Our work differentiates itself from these approaches by introducing a general and efficient algorithm for searching BB codes in the form described in Eq. (5), which is also the form of codes used by related works. Furthermore, we proposed a novel algorithm that allows us to search for a subclass of BB codes with the desired dimensions, which we call coprime-BB codes. Coprime-BB codes generalize the form in Eq. (5) by allowing mixed terms but also have a restriction of using coprime circulant matrix sizes, making them unattainable through searches limited to the polynomial form of this equation. Thus, coprime-BB codes provide more flexibility for different scenarios on top of the originally proposed BB codes.
@@ 3.1 Code equivalence
share the same parameters, allowing us to search within only one class of these codes. These equivalences can also be viewed as special cases of Theorem 6 in [15].
@@ 4 A Novel Layout Optimization for Coprime-BB Codes
Discussion: In the CBB layout, qubits are arranged into long stripes forming a $1\times lm$ grid, rather than a $l\times m$ rectangular array. This geometry places increased demands on acousto-optic deflectors, since the larger number of columns requires a correspondingly higher number of resolvable radio-frequency tones, which may constrain the maximum achievable array size due to finite AOD bandwidth and spot-resolution limits. However, coprime-BB codes are known to be asymptotically limited [18], and are therefore not intended for very large code distances. As a result, our focus is on short-to-medium code lengths, for which the required number of sites per row remains moderate. Recent experimental demonstrations of neutral-atom arrays have achieved grids as large as $120\times 12$ [19] suggesting that resolving on the order of $10^{2}$ atoms per row is feasible with current technology.
In addition, ancilla transport in the CBB layout typically involves longer distances per move than in BB layouts. Although our analysis indicates that the total number of movement operations, as well as the overall transport time for the codes considered here, is reduced, this reduction is achieved at the cost of higher average transport velocities. As discussed in [20, 21], the start-stop jerking effect could be a major source of heating and atom loss. The trade-off between fewer start–stop events and increased velocity makes the net impact on transport-induced noise, such as motional heating, difficult to assess without detailed experimental characterization. In summary, the CBB layout is more suitable for short-to-medium code lengths on near-term devices.
@@ 6 Related Work
Recently, various constructions based on qLDPC codes have been proposed to achieve different objectives. Koukoulekidis et al. [24] proposed algebraic extensions to expand a small GB code into a family of larger GB codes by selecting a sequence of expansion factors. The authors also introduced scalable codes that embed the original short codes into extended codes, enhancing scalability in superconducting architectures. Voss et al. [25] expanded the concept of BB codes by introducing an additional type of indeterminate variable, leading to the creation of trivariate bicycle (TB) codes. These new codes reduce the weight of stabilizers from 6 to 4-5, making them more practical for hardware implementation. However, it is important to note that some of these codes exhibit a lower rate or distance compared to BB codes of similar length. Eberhardt et al. [16] investigated the algebraic structure of BB codes and uncovered certain symmetry properties. These properties allow for the explicit construction of logical operators and certain fault-tolerant gates for BB codes. Shaw et al. [26] proposed a “morphing” circuit design for syndrome extraction on BB codes. The proposed circuit design has only six rounds of CNOT gates instead of seven [5]. By applying the proposed circuit, the authors discovered a new family of BB codes, including codes that have the same $\llbracket n,k,d\rrbracket$ parameters as [5]. The authors also provide a sufficient condition for the circuit to be applied to the other two-block group algebra (2BGA) [15] codes.